OpenStudy (anonymous):

Last problem of the day! Find f. f ''(t) = 2e^t + 3 sin t, f(0) = 0, f(π) = 0

OpenStudy (callisto):

f ''(t) = 2e^t + 3 sin t f'(t) = ∫f ''(t)dt = ∫ 2e^t + 3 sin t dt = 2e^t - 3 cos t + C1 f(t) = ∫f '(t)dt = ∫(2e^t - 3 cos t + C1)dt = 2e^t - 3 sin t + (C1)t +C2 f(0) = 2e^(0) - 3 sin (0) + (C1)(0) +C2 = 0 2 -0 +0 +C2 =0 C2 = -2 f(π) = 0 2e^(π) - 3 sin (π) + (C1)(π) +C2 = 0 2e^(π) -0 + (C1)(π) -2 =0 C1 = [2-2e^(π)] /π therefore f(t)= 2e^t - 3 sin t + { [2-2e^(π)] t} /π -2